# Neural Attention Search

### v2

@@ 3.3 Efficient Inference with Different Token Types
Similar to the vanilla transformer forward process, when new tokens arrive, we concatenate them with the existing KV caches, generating new masks and computing the attention output. After that, we start our post-update process: we first check the state of each token to decide if we could remove them or keep them in the cached values. Since different heads might disagree on the types of the same token, we record both the sizes for Global Tokens ($Size_{G}$) and Local Tokens ($Size_{L}$) for all the heads. New Sliding Window Tokens do not change these sizes since they will always be placed separately. However, when a new Global Token for any head arrives, we remove all the Local Tokens from the corresponding heads and place the new Global Token right after the existing Global Tokens and then update our $Size_{G}$ and $Size_{L}$ accordingly. The same strategy is applied when new $LocalTokens$ arrive: we place them at the end of the Local Tokens and increase the number for Local Tokens.
@@
Figure 5 illustrates an exemplary case to update the KV caches in NAtS with a sliding window size of 3. The first two places store the most recent tokens and use a mask to mask out the non-Sliding Window Tokens (yellow tokens). Since the Token 8 for Head 1 is Sliding Window Token and Global Token for Head 2. We first move both tokens to the beginning of our cache to replace the old one. After that, we drop Token 8 in Head 1 since it has already moved to the sliding window caches. Then, since Token 8 in Head 2 is a Global Token, we drop all local tokens after the last Global Token (Token 1). Hence, Tokens 4 and 6 are removed from the cache, and we place Token 8 after Token 1. After that, the $Size_{G}$ is updated from $[5,3]$ to $[5,4]$ and the $Size_{L}$ is updated from $[0,2]$ to $[0,0]$. Since both new tokens are merged into the existing tokens, we do not need the extra space to store these new tokens.
@@ 4 Experiments
We implement NAtS based on the Flash Attention 2 (Dao, 2024) implementation on triton. All the operations that we proposed in Section 3.2 have at most $O(\mathit{L})$ complexity. In our experiments, we first train NAtS parameters jointly within a transformer model from scratch. We then apply NAtS to fine-tune a large language model.
@@ 4.1 Training a Transformer From Scratch
We first apply NAtS  to train a GPT2 small style (Radford et al., 2021) transformer model from scratch. Following the setting from NanoGPT (Karpathy, 2022), this model has 128M parameters with 12 layers and 12 heads with a hidden dimension of 768. Instead of the learnable position encoding, we apply rotary embeddings (Su et al., 2024) to each transformer layer. We train this model on the PG-19 Language Modeling Benchmark (Rae et al., 2020). This dataset contains books extracted from Project Gutenberg (PG, 19) with about 2B tokens in the training sets. We train all models with a context length of 1024 and batch size of 480 (using gradient accumulation). We train all models for 120 000120000120\,000120 000 iterations on the training sets on a computation node with four Nvidia H100 GPUs. We evaluate all models on the test sets of PG19 with a context length of 1024.
As a baseline, we train another dense transformer model under the same hyperparameter setting. During inference time, we compare NAtS  with the following baselines besides the full Transformer: (i) Streaming LLM (Xiao et al., 2023) only preserves the first few starting and the most recent few tokens for future prediction. (ii) H2O (Zhang et al., 2023b) first computes the attention map and only preserves the tokens with the top-k attention scores. H2O and Streaming LLM are training-free approaches that control the sparsity with pre-defined hyperparameters during inference time. H2O needs to define the recent sliding window size and the number of Heavy Hitter (HH) tokens; Streaming LLM requires the number of tokens at the beginning of the sequence (Sink Tokens) and sliding window size. In contrast, NAtS  controls this sparsity by the sparsity hyperparameter value $\lambda$. Hence, we train multiple models with different $\lambda$. However, in our experiments, we observe that the attention map sparsity values converge much faster than the model loss. We could quickly estimate the attention map sparsity within the first 10,000 iterations to check if this sparsity value satisfies the user requirement and early-stop the runs that do not satisfy the requirements (Jamieson & Talwalkar, 2016).
@@
We provide different hyperparameters for H2O and Streaming LLM (with the sliding window size of 32, 64, 128, and 256, plus the same number of HH/ Sink tokens). This results in a corresponding KV budget fraction ranging from 6.25% to 50%. Meanwhile, we train NAtS with the following $\lambda$: $70,1e-9,5e-9,1e-8,1e-$ 7.
@@
Figure 6 shows the perplexity of the PG19 test sets from different model settings. The x-axis indicates the fraction of KV caches applied to generate the last token in the input sequence. As expected, larger $\lambda$ results in a smaller KV cache fraction. Interestingly, even if $\lambda$ is set as 00, the KV cache budget can still be reduced to around 60% with slightly lower perplexity. This might indicate that the model also tends to sparsify the input information to focus more on the relevant parts from the input data (Ye et al., 2024).
@@
Under the budget of around 50%, all the approaches perform similarly to the full transformer while NAtS has a relatively lower perplexity and requires less budget. However, as the available budget decreases, NAtS’s perplexity remains nearly the same until a KV cache size of around $10\%$. In contrast, H2O and Streaming LLM start to degenerate their performance starting from a $25\%$ budget allocation. The most sparse model in NAtS family only contains roughly $3\%$ of the KV cache size (meaning roughly 30 cached tokens per layer on average) and achieves a lower perplexity compared to H2O with $12.5\%$ of the KV caches and Streaming LLM with $25\%$ of the KV caches.
4.2 Fine Tune a Large Language Models
We now apply NAtS to fine-tune an existing large language model (Grattafiori et al., 2024; Jiang et al., 2023) in the long context scenario. We evaluate the fine-tuned models on the Needle-in-a-Haystack (NIAH) test (Kamradt, 2022) and LongBench dataset (Bai et al., 2024). NIAH places a random fact or statement (needle) in the middle of a long context and asks the model to retrieve it from the context. LongBench evaluates the model’s capacity to understand the information in different long context scenarios, such as single-document and multi-document QA, summarization, few-show learning, synthetic tasks, and code completion.
In addition to the baselines in Section  4.1, we use DuoAttention (Xiao et al., 2024) as another baseline. DuoAttention evaluates the importance of each attention head and only assigns full attention budgets to the heads with high scores (the so-called retrieval heads) while applying streaming to the other unimportant heads. DuoAttention shares the same idea as NAtS, where the importance of different KV caches should be learned instead of the pre-defined rules. However, instead of learning the importance of the head level, we aim at learning this information directly on the token level.
@@ 4.2 Fine Tune a Large Language Models
To allow our models to learn to evaluate the role of each token within different contexts, we construct a new training dataset that follows the construction rules of LongBench. Some of LongBench’s tasks are collected from the test sets of the previous benchmarks. Hence, we first collect the training sets from these benchmarks and construct these datasets following the data structure in LongBench. Overall, this dataset contains 6 43664366\,4366 436 pieces of data with a maximum context length of 16 0001600016\,00016 000. Further details on the dataset collection can be found in the appendix.
@@
We only fine-tune the Attention Score Layer while keeping all the other parameters in the network fixed. Hence, we aim to approximate the original output from the corresponding base LLM. This approach is similar to DuoAttention (Xiao et al., 2024). However, since we want the model to capture the overall context information, we update Attention Score Layer with all the output from the full attention layer:
@@
where $\mathit{B}$ is the Batch Size. Additionally, we control the sparsity with the regularization value $\lambda$ instead of the additional loss item defined in DuoAttention, and therefore, only optimize NAtS with $\mathcal{L}_{distill}$ as the loss function. We fine-tune two long-context models (Llama-3-8B-Instruct-Gradient-1048k (Llama-3 8B Gradient Instruct 1048k, ) and Mistral-7B-v0.3-Instruct  (Jiang et al., 2023)) on two Nvidia H100 GPUs for one epoch using AdamW (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with a learning rate of 3e-4 with a warm-up from 3e-5 in the first 20% steps and reducing back to 3e-5 in the last 20% steps. We apply different $\lambda$ to allow for different sparsity. For a fair comparison, we set the sliding window size W as 256, the same as DuoAttention.
@@
Since we only optimize the Attention Score Layer, the number of learnable parameters is $n_{layers}\cdot d_{model}\cdot n_{heads}\cdot{n_{options}}$. Where $n_{options}$ is the number of options in our search space (in this case, it is $3$). Hence, the size of the parameters that need to be stored is neglectable (it only takes roughly $13$MB on disk for each set of Attention Score Layer) compared to the weights of the LLM. Hence, users could pre-collect all these weights and apply the one that fits their compression requirement.
For the Needle In a Haystack task, we test all models with a context length ranging from 1 00010001\,0001 000 to 32 0003200032\,00032 000 with the needle positions ranging from 00 to $1$. We assign a $25\%$ KV cache budget to each of the baseline approaches. For NAtS, we choose the model with the compression rate closest to $25\%$, which corresponds to a model trained with $\lambda=1e-7$ and an average KV budget size of $15\%$.
The results on Needle in a Haystack test are shown in Figure LABEL:fig:expniah. DuoAttention, H2O, and Streaming LLM all maintain a sliding window with a fixed size. Therefore, they could well capture the needles inserted at the text’s tail. However, all the baselines struggle with the correct answers as the needle goes into the middle of the context. In contrast,  NAtS efficiently recognizes the necessary tokens across different layers, heads and positions to preserve only tokens that contain the important information.
We also plot the KV budget percentage on the bottom right of Figure LABEL:fig:expniah. Since we set our Sliding window size W as 256, the KV budget for the first 1000 tokens is roughly $30\%$. However, as the context length increases, the compression rates gradually decrease to the over compression rate of $15\%$. This shows NAtS could save the important information uniformly across the context in different positions.
Full Duo SLLM H2O NAtS
NarrativeQA 27.3 17.4 22.1 25.75 23.09(13%)
Qasper 28.9 16.4 13.0 16.9 29.01(26%)
MultiFieldQA-en 52.3 35.8 40.6 29.1 48.72(22%)
MultiFieldQA-zh 50.9 37.9 34.4 28.2 46.03(27%)
HotpotQA 40.6 26.8 28.1 32.8 41.7(17%)
2WikiQA 29.3 22.2 23.2 24.4 27.39(20%)
Musique 24.4 12.0 17.9 18.1 21.0(16%)
DuReader (zh) 30.2 26.3 31.3 17.7 28.44(20%)
GovReport 34.1 22.4 28.8 24.1 33.47(20%)
QMSum 24.5 20.9 21.5 21.4 23.88(15%)
MultiNews 27.7 24.0 24.4 22.4 27.77(32%)
VCSUM (zh) 11.3 4.8 10.6 13.14 9.4(19%)
TREC 71.0 53.5 61.5 55.5 68.5(22%)
TriviaQA 87.7 77.8 82.8 84.9 87.42(18%)
SAMSum 42.5 41.0 39.4 38.8 41.78(14%)
LSHT 38.0 21.5 25.0 22.2 35.5(20%)
Passage Count 1.0 2.09 2.0 2.0 1.59(23%)
PassageRetrieval-en 80.5 46.5 25.5 47.0 67.0(15%)
PassageRetrieval-zh 61.2 29.4 14.0 34.2 41.07(24%)
LCC 37.5 36.2 43.5 44.74 40.11(35%)
RepoBench-P 38.0 39.1 42.0 41.9 42.63(25%)
@@
Table 1: LongBench Results with 25% Budget Allocation for full attention (Full), DuoAttention (Duo), Streaming LLM (SLLM), H2O and NAtS on the LLama8B model. The numbers in the brackets for NAtS is the used value KV cache sizes. We bold the models with the best performance besides the base Full Attention Models .
@@
For the LongBench tasks, we evaluate all the baselines with the $50\%$ and $25\%$ of the KV cache sizes. We show the result with KV cache size of $25\%$ for LLama3-8B in Table 1. The model that is used in this task is a model with $\lambda=1e-7$. NAtS achieves the best performance on most of the dataset with (in many cases) smaller KV cache values. We provide further results, including results with Mistral-7B and  NAtS trained with other parameters, in the appendix.
4.2.1 KV Size Distribution
Figure LABEL:fig:KVSizeDistributions shows the distribution of different KV budgets on the narrativeQA (Kociský et al., 2018) and RepoBench-p (Liu et al., 2024) dataset. The overall trends are similar: The first few layers and some specific heads require more layers than the other heads. However, as the model goes deeper, the model tends to drop more tokens for the NarrativeQA dataset while still maintaining many tokens for RepoBench-P, especially after layer 16. This shows that a fixed rule might not always fit all the sequences and there is a need to adjust the sampling strategy with different inputs.

### v3

@@ 3.3 Efficient Inference with Different Token Types
Similar to the vanilla transformer forward process, when new tokens arrive, we concatenate them with the existing KV caches, generating new masks and computing the attention output. After that, we start our post-update process: we first check the state of each token to decide if we can remove them or keep them in the cached values. Since different heads might disagree on the types of the same token, we record both the sizes for Global Tokens ($Size_{G}$) and Local Tokens ($Size_{L}$) for all the heads. New Sliding Window Tokens do not change these sizes since they will always be placed separately. However, when a new Global Token for any head arrives, we remove all the Local Tokens from the corresponding heads and place the new Global Token right after the existing Global Tokens and then update our $Size_{G}$ and $Size_{L}$ accordingly. The same strategy is applied when new $LocalTokens$ arrive: we place them at the end of the Local Tokens and increase the number for Local Tokens.
@@
Figure 5 illustrates an exemplary case to update the KV caches in NAtS with a sliding window size of 3. The first two places store the most recent tokens and use a mask to mask out the non-Sliding Window Tokens (yellow tokens). Since Token 8 for Head 1 is Sliding Window Token and Global Token for Head 2. We first move both tokens to the beginning of our cache to replace the old one. After that, we drop Token 8 in Head 1 since it has already moved to the sliding window caches. Then, since Token 8 in Head 2 is a Global Token, we drop all local tokens after the last Global Token (Token 1). Hence, we place Token 8 after Token 1 and remove Tokens 4 and 6. The $Size_{G}$ is thenupdated from $[5,3]$ to $[5,4]$ and the $Size_{L}$ is updated from $[0,2]$ to $[0,0]$: both new tokens are merged into the existing tokens, we do not need the extra space to store them.
@@ 4 Experiments
We implement NAtS based on the Flash Attention 2 [17] implementation on triton. In our experiments, we first train NAtS parameters jointly within a transformer model from scratch. We then apply NAtS to fine-tune a large language model to show that NAtS could efficiently reduce the KV cache size required while maintaining the model performance. The codes for  NAtS implementation and experimental designs can be found under: https://github.com/automl/NeuralAttentionSearch
@@ 4.1 Training a Transformer From Scratch
We first apply NAtS to train a GPT2 small style [67] transformer model with 128M parameters from scratch on the PG-19 Language Modeling Benchmark [68] with four Nvidia H100-PCIe GPUs and evaluate them on the test sets of PG19 with a context length of 1024. Further details can be found under the appendix  E.1.
As a baseline, we train another dense transformer model under the same hyperparameter setting. During inference time, we compare NAtS with the following baselines besides the full Transformer: (i) Streaming LLM [82] only preserves the first few starting and the most recent few tokens for future prediction. (ii) H2O [93] first computes the attention map and only preserves the tokens with the top-k attention scores. H2O and Streaming LLM are training-free approaches that control the sparsity with pre-defined hyperparameters during inference time. In contrast, NAtS controls this sparsity by the sparsity hyperparameter value $\lambda$. Hence, we train multiple models with different $\lambda$. However, in our experiments, we observe that the attention map sparsity values converge much faster than the model loss. We could quickly estimate the attention map sparsity within the first 10,000 iterations to check if this sparsity value satisfies the required sparsity and early-stop the runs that do not satisfy the requirements [38].
@@
We provide different hyperparameters for H2O and Streaming LLM for different sparsity (with the sliding window size of 32, 64, 128, 256 plus the same number of HH for H2O and 64, 128, 256, 512 plus 8 Sink tokens for Streaming LLM). This results in a corresponding KV budget fraction ranging from 6.25% to 50%. Meanwhile, we train NAtS with the following $\lambda$: 0,1e9,5e9,1e8,1e70,1e-9,5e-9,1e-8,1e-7.
@@
Figure 6 shows the perplexity of the PG19 test sets from different model settings. The x-axis indicates the fraction of KV caches applied to generate the last token in the input sequence. As expected, larger $\lambda$ results in a smaller KV cache fraction. Interestingly, even if $\lambda$ is set as 0, the KV cache budget can still be reduced to around 35% with slightly higher perplexity compared to the full attention. This might indicate that the model also tends to sparsify the input information to focus more on the relevant parts from the input data [87].
@@
H2O maintains nearly lossless perplexity with a budget of 50%. However, as the available budget decreases,  NAtS’s perplexity remains nearly the same until a KV cache size of around 10%. In contrast, H2O and Streaming LLM start to degenerate their performance starting from a $25\%$ budget allocation. The most sparse model in NAtS family only contains roughly $3\%$ of the KV cache size (meaning roughly 30 cached tokens per layer on average) and achieves a lower perplexity compared to H2O with $12.5\%$ of the KV caches.
4.2 Fine Tune a Large Language Model
@@
We now apply NAtS to fine-tune an existing large language model [32, 40] in the long context scenario. We construct a new training dataset that follows the construction rules of LongBench as a training set to fine-tune LLM models with NAtS. Some of LongBench’s tasks are collected from the test sets of the previous benchmarks. Hence, we first collect the training sets from these benchmarks and construct these datasets following the data structure in LongBench. Additionally, we also add the passkey-retrieval dataset introduced in DuoAttention [81]. Overall, this dataset contains 7 0007\,000 instances with a maximal context length of 16 00016\,000. Further details can be found in Appendix E.2.
@@
We only fine-tune the Attention Score Layer while keeping all the other parameters in the network fixed to approximate the original output from the corresponding base LLM. This approach is similar to DuoAttention [83]. However, since we want the model to capture the overall context information, we update Attention Score Layer with all the output from the full attention layer:
@@
where $\mathit{B}$ is the Batch Size. We only use the context information from the real-world dataset to ensure that the token information does not rely on specific prompts. However, the synthetic dataset might contain duplicated contexts, and hence, we only optimize the synthetic dataset with its corresponding labels [83]. Additionally, we control the sparsity with the regularization value $\lambda$ instead of the additional loss item defined in DuoAttention, and therefore, only optimize NAtS with $\mathcal{L}_{distill}$ as the loss function. We fine-tune two long-context models (Llama-3.1-8B-Instruct [32] and Mistral-7B-v0.3-Instruct [40]) on two Nvidia H100 PCIe GPUs for one epoch using AdamW [45, 57] with a learning rate of 2e-3 with a warm-up from 2e-4 in the first 20% steps and reducing back to 2e-4 in the last 20% steps. We apply different $\lambda$ to allow for different sparsity. For a fair comparison, we set the sliding window size W as 256, the same as DuoAttention.
In addition to the baselines in Section 4.1, we also evaluate other state-of-the-art KV eviction strategies, including SnapKV [52], AdaKV [25], ChunkKV [55], PyramidKV [14], CriticalKV [26]. These approaches rely on a sliding window from the query matrix during the pre-filling stage to identify the important tokens. Unlike NAtS and the other approaches listed above that update the KV caches on the fly, these approaches only evict the KV cache once after the prefilling stage. Additionally, we evaluate DuoAttention [83] and MoA [27]. Both approaches extend the existing streaming LLMs approach to provide different budgets for different attention heads. These two methods share the same idea as NAtS, where the importance of different KV caches should be learned instead of the pre-defined rules. However, instead of learning the importance of the head level, we aim at learning this information directly on the token level.
@@
Since we only optimize the Attention Score Layer, the number of learnable parameters is $n_{layers}\cdot d_{model}\cdot n_{heads}\cdot{n_{options}}$, where $n_{options}$ is the number of options in our search space (in our case, it is $3$). Hence, the size of the parameters that need to be stored is negligible (i.e., it only takes roughly 1313MB on disk for each set of Attention Score Layer) compared to the weights of the LLM. Hence, users could pre-collect all these weights and apply the one that fits their compression requirement.
We evaluate all the baseline results on the RULER [36] and LongBench [11]. RULER is a synthetic benchmark that evaluates a model’s long-context capabilities across 4 task categories. LongBench evaluates the model’s capacity to understand the information in different long context scenarios. Following the setting from KV Press [63], we ask all the one-time KV eviction approaches to evict the KV cache once the model receives all the context information before the question-related information arrives. For the other approaches, we evict the KV caches dynamically when new tokens are generated.
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
4k 95.06 94.25 59.37 83.03 54.90 70.03 69.19 87.52 56.41 44.63 95.32 (51%)
8k 93.22 92.42 48.58 81.51 54.10 71.43 67.20 89.01 54.55 19.80 93.30 (45%)
16k 90.53 88.24 50.97 65.55 54.38 76.32 67.07 88.03 58.17 15.16 90.01 (42%)
32k 85.95 84.65 45.68 11.48 47.13 72.48 55.88 82.88 48.05 12.79 86.16 (41%)
64k 84.01 83.54 51.40 0.07 69.61 78.10 71.05 81.93 70.49 10.86 84.09 (41%)
128k 73.83 73.11 46.95 0.00 39.26 64.49 41.61 70.53 40.02 11.05 74.77 (44%)
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
4k 95.06 73.18 37.30 71.95 35.67 44.34 51.59 57.73 35.60 35.60 93.93 (24%)
8k 93.22 62.83 28.29 65.34 33.92 41.46 45.87 62.83 35.17 25.69 90.29 (19%)
16k 90.53 52.01 28.19 47.82 32.74 41.61 49.40 67.08 34.35 23.94 88.87 (16%)
32k 85.95 53.50 30.54 9.71 32.17 40.24 39.19 64.41 32.49 20.17 84.40 (15%)
64k 84.01 65.57 31.42 0.04 57.36 65.98 57.01 70.85 58.02 19.29 79.36 (15%)
128k 73.83 53.05 31.41 0.00 27.36 34.33 37.35 56.51 27.34 13.40 64.26 (16%)
@@
Table 1: Ruler results with 50% KV budgets (top) and 25% KV budget (bottom) on LLama 3.1. All Full models use the 100% KV budgets. We mark the actual KV budgets used by NAtS in the bracket. Models with the best performance besides the base Full Attention Models are bold
The results on the Ruler Benchmark with LLama 3.1 are shown in Table  1. NAtS uses $\lambda=3e-7$ for the 25% KV size level and $\lambda=5e-8$ for the 50% KV size level. NAtS achieves nearly lossless scores across different input context lengths. When the KV budget drops to 25%, NAtS achieves 96% of the full attention scores with much smaller KV budget sizes, while the other baselines could no longer keep their performance with the decreased KV budgets. Results with Mistral-7B-Instruct-v0.3 can be found under Appendix F.
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
NarrativeQA 31.35 22.62 27.83 21.07 27.08 29.02 26.27 30.48 28.02 19.19 30.53(13%)
Qasper 24.73 18.99 14.53 15.00 14.24 14.84 12.71 15.79 13.43 18.57 23.76(27%)
MultiFieldQA-en 29.46 25.36 16.47 18.24 17.72 19.31 17.44 21.74 17.72 14.75 28.94(24%)
MultiFieldQA-zh 60.01 54.95 33.94 29.14 33.77 35.34 33.49 36.30 32.76 21.23 61.18(22%)
HotpotQA 17.06 16.06 13.95 16.47 15.58 16.12 16.50 16.73 14.23 10.18 16.06(20%)
2WikiQA 16.64 16.18 13.53 11.28 13.45 13.17 13.41 13.89 12.36 11.42 17.05(23%)
Musique 11.59 8.41 8.96 11.13 10.81 10.68 11.05 11.25 9.40 5.93 11.42(18%)
DuReader (zh) 35.56 32.18 18.64 24.77 23.57 24.57 24.24 24.68 23.56 22.33 34.28(17%)
GovReport 34.30 27.63 28.20 26.90 28.51 28.28 28.86 29.58 27.53 25.24 33.82(19%)
QMSum 23.30 22.28 20.67 20.39 21.41 22.05 21.27 22.27 21.98 19.59 23.11(17%)
MultiNews 27.13 24.58 21.76 22.80 23.14 23.50 22.72 24.03 23.00 23.30 26.72(33%)
VCSUM (zh) 16.36 14.97 14.26 14.85 14.80 15.18 14.92 15.46 15.21 14.58 15.61(14%)
TREC 72.50 58.50 66.00 55.00 53.50 59.00 54.50 62.00 48.50 60.50 72.00(26%)
TriviaQA 91.15 82.81 90.69 89.98 91.00 91.55 90.72 90.97 90.92 71.51 91.61(22%)
SAMSum 43.72 40.14 41.97 41.98 43.59 43.11 42.32 43.79 43.15 42.30 44.21(16%)
LSHT 46.50 35.00 35.50 27.50 45.00 44.00 43.50 46.00 45.00 22.00 47.50(16%)
Passage Count 6.63 4.50 7.22 3.40 5.52 5.14 4.11 5.90 4.70 3.79 7.12(20%)
PassageRetrieval-en 97.98 91.65 94.81 72.25 90.36 90.70 90.38 94.70 88.32 28.21 95.81(19%)
PassageRetrieval-zh 77.99 55.22 28.97 52.13 73.06 78.74 55.95 80.15 74.66 24.78 78.5(20%)
LCC 54.10 54.27 52.38 54.52 55.40 54.93 54.12 54.26 54.90 54.14 53.54(38%)
RepoBench-P 51.39 57.26 49.65 55.54 51.80 52.44 52.87 52.57 52.41 53.35 53.48(28%)
Table 2: LongBench Results with 25% Budget Allocation. The Full model uses the 100% KV budgets. The numbers in the brackets for NAtS is the used value KV cache sizes. We bold the models with the best performance besides the base Full Attention Models .
@@
For the LongBench tasks, we evaluate all the baselines with $50\%$ and $25\%$ of the KV cache sizes. We show the result with KV cache size of $25\%$ for LLama3.1-8B in Table 2. The model that is used in this task is a model with $\lambda=3e-7$. NAtS achieves the best performance on most of the datasets with (in many cases) smaller KV cache values. We provide further results, including results with Mistral-7B and  NAtS trained with other parameters, in Appendix  F.
4.3 Latency Evaluation
[…]
Figure 7: Memory and latency usage during pre-filling (left) and decoding (right)